Find the voltage across a resistor if the current flowing through it is 5A. And also determine the power and 

energy consumed by the resistor if the current is allowed to pass through the resistor for 20 min.


1
Expert's answer
2021-04-19T17:09:59-0400

According to the Ohm's law, the voltage accross the resistor is given as follows:


"U = IR"

where "R" is the resistance of the resistor, and "I = 5A" is the current that passes throgh it. Since we are not given with the resistance, we can not calculate the voltage.

The power is given as follows:


"P = UI = I^2R"

and the energy consumed:


"Q = Pt = I^2Rt"

where "t = 20min = 1200s" is the time.


Answer. "U = IR, \space P = I^2R, \space Q = I^2Rt".